# Earthfarseer: Versatile Spatio-Temporal Dynamical Systems Modeling in One Model

### v1

@@ Introduction
Modeling spatio-temporal (ST) physical dynamics involves estimating states and physical parameters from a sequence of observations (Benacerraf 1973; Newell 1980). Generally, the understanding of a physical process is based on plenty of physical laws, such as Newton’s second law (Pierson 1993) and Conservation of energy law (Sharan et al. 1996; Egan and Mahoney 1972; Shan, Yuan, and Chen 2006; Wang et al. 2022b). As tailor-made techniques, dynamical systems, primarily rooted in diverse physical systems, have been demonstrated to conform to most fundamental principles of real-world physical phenomena (Chmiela et al. 2017; Hoedt et al. 2021; Greydanus, Dzamba, and Yosinski 2019), where such phenomena can be mostly recognized by existing mathematical frameworks. To this end, the modeling of dynamic systems has increasingly become a generic approach that yields numerous versatile techniques for various applications (Wiggins, Wiggins, and Golubitsky 2003; Hale and Koçak 2012; Humar 2012; Wang, Zhou, and Jin 2020). Actually, such dynamic systems can naturally model the time-varying evolution including both intricate natural system like meterology dynamics (Wiggins, Wiggins, and Golubitsky 2003; Hale and Koçak 2012; Humar 2012; Harish and Kumar 2016) and complex society system of human mobility like traffic evolutions (Ji et al. 2022; Chen et al. 2022a; Vinuesa and Brunton 2022; Yang et al. 2022).
@@
Intriguingly, the fascination of machine learning with physical phenomena has significantly intensified in recent years (Shi et al. 2015; Long et al. 2018; Greydanus, Dzamba, and Yosinski 2019; Donà et al. 2020). Many research efforts have turned to video understanding (Chang et al. 2021; Wang et al. 2022c; Chang et al. 2022; Gao et al. 2022b) and physical-guided (Jia et al. 2021; Aoyagi, Murata, and Sakaino 2021; Lu et al. 2021) deep learning to capture ST characteristics in a data-driven manner. These approaches usually design various spatial or temporal components tailored for effectively characterizing a specific scenario or a dedicated task. Although many ST frameworks deliver higher accuracy than the simple ones, they inevitably suffer from partial drawbacks ($\cal D$) outlined below:
  • ${\cal D}1$:
    Lack of local fidelity. Given the inconsistency and distinctive local-global dependencies in dynamic systems, existing modules usually focus on global regularity but fail to preserve the local fidelity – local dynamics may differ from global dynamics (See example in Fig 1 and Fig 8 in Appendix A).
  • ${\cal D}2$:
    Poor predictions over long time-steps. Complex and continuous dynamic systems often exhibit intricate temporal correlations, leading to poor performance in long-term predictions (Isomura and Toyoizumi 2021).
  • ${\cal D}3$:
    Low scalability. The intricate and convoluted component designs confine the model only capable to resolve specific tasks (e.g., super-resolution (Liang et al. 2019), flow prediction (Pan et al. 2019b)) or datasets, leading to limited scalability.
  • ${\cal D}4$:
    Inefficiency. The well-designed but cumbersome ST blocks not only contribute to the inefficient training (Wang et al. 2023, 2022a) process but also pose challenges to the model deployment.
@@
For temporal correlations, we design a temporal dynamic evolution module, TeDev, which effectively captures the continuous dynamic evolution within low-dimension space. In comparison to traditional modeling of discrete static frames (Shi et al. 2015; Walker, Razavi, and Oord 2021; Wang et al. 2022c), our model undergoes a transformation from the continuous time domain to the frequency domain through Fourier transformation, better preserving the long-term dependence of spatio-temporal data. Further, different from the prevailing neural ordinary differential equation (ODE) algorithms (Park et al. 2021), our model showcases the great prominence in efficiency and long-term prediction tasks when captures intricated dynamics of systems without differential equation based nonlinear features. Through the exploitation of a low-parameter linear convolution projection, TeDev can effciently and accurately predict arbitrary future frames (right side of Fig 2).
@@
Our work shares common ground with several lines of research, and we will now summarize these studies below.
Spatio-temporal prediction methods can be roughly divided into CNN-based (Oh et al. 2015; Mathieu, Couprie, and LeCun 2015; Tulyakov et al. 2018), RNN-based (Ranzato et al. 2014; Srivastava, Mansimov, and Salakhudinov 2015; Villegas et al. 2017, 2018; Kim, Ahn, and Bengio 2019; Wang et al. 2022c), and other models including the combinations (Weissenborn, Täckström, and Uszkoreit 2019; Kumar et al. 2019) and transformer based models (Dosovitskiy et al. 2020; Gao et al. 2022a; Bai et al. 2022). While there are several existing models based on graph neural networks (GNNs), their primary focus is on handling graph data (Sun et al. 2020; Wang, Cao, and Philip 2020; Jiang et al. 2021; Jiang, Ji, and Li 2019; Wang et al. 2022b), which go out of the scope of our work.
Video Prediction has become a crucial research topic in the multimedia community, resulting in the proposal of numerous methods to tackle this challenge. Early studies primarily focused on analyzing spatio-temporal signals extracted from RGB frames (Shi et al. 2015; Lotter, Kreiman, and Cox 2017; Wang et al. 2022c). Recently, there has been a growing interest in integrating video prediction with external information such as optical flow, semantic maps, and human posture data (Liu et al. 2017; Lee et al. 2021; Wang et al. 2018; Pan et al. 2019a). However, in real-world applications, accessing such external information may not always be feasible (Pan et al. 2019a; Wu et al. 2020; Hu et al. 2023). Moreover, the current solutions still exhibit suboptimal efficiency and effectiveness when dealing with high-resolution videos. In this study, we concentrate on modeling continuous physical observations, which can sometimes be interpreted as a video prediction task.
@@
The dynamical systems dictate the principles that govern the transition of the current system state to its future continuous state. Formally, given an environment space $X$ and a state space $S\subset X$, a dynamical system can be represented by a function $F$:
@@
As shown in Fig 3(b), stream morph operator first converts the discrete ST sequence into a continuous stream, followed by the MSFC module and the Fourier-based computational unit to complete the hidden feature extraction. Stream morph combines the channel dimension and the time dimension, enabling the transformation of discrete temporal dynamics into a continuous and irregular representation. The MSFC module stands out from traditional methods by utilizing a multi-scale fully convolutional (MSFC) architecture with various kernel sizes. This architecture enables feature extraction from a global to local perspective, with larger kernel sizes (7 and 11) capturing global information distribution and smaller sizes (1 and 3) facilitating locality. TeDev also integrates FFT/IFFT module, which converts time-domain signals into frequency-domain signals (See Eq $4\sim 6$), facilitating a better understanding and analysis of signal components. As a result, TeDev captures information at different time scales and avoids losing temporal information during spatiotemporal prediction. Formally, TeDev can be described as follows:
@@
In this section, we empirically demonstrate the superiority of our framework on seven datasets, including two human social dynamics system (TaxiBJ+ (Zhang, Zheng, and Qi 2017), KTH (Schuldt, Laptev, and Caputo 2004)), five natural Scene dynamical systems (SEVIR (Veillette, Samsi, and Mattioli 2020), RainNet (Chen et al. 2022b), Pollutant-Diffusion (PD), Reaction-Diffusion (Chen et al. 2022a)) and 2D shallow water Equations (2DSWE) (Takamoto et al. 2022), and a synthetic systems (MovingMNIST (Srivastava, Mansimov, and Salakhudinov 2015)). In the subsequent section, we will provide a detailed introduction to the dataset and baseline, along with the corresponding experimental settings and results.
@@
  • ${{\cal B}_{1}}$.
    Video Prediction Models: We select ConvLSTM (Shi et al. 2015), PredRNN-v2 (Wang et al. 2022c), E3D-LSTM (Wang et al. 2019) and SimVP-v2 (Tan et al. 2022) as some of the most representative and advanced RNN architecture models in recent years.
  • ${{\cal B}_{2}}$.
    Spatio-temporal Series Modeling: We conduct experiments on advanced Transformer architecture models, including Vision Transformer (Dosovitskiy et al. 2020), Swin Transformer (Liu et al. 2021), Rainformer (Bai et al. 2022) and Earthformer (Gao et al. 2022a).
  • ${{\cal B}_{3}}$.
    Physics-guided Neural Networks: We use modeling methods that incorporate PDE or ODE in the model as baseline models for comparison, including Ad-fusion model, PhyDnet (Guen and Thome 2020), Vid-ODE (Park et al. 2021), PDE-STD (Donà et al. 2020), and FourcastNet (Pathak et al. 2022), and FourcastNet represents a meteorological modeling model in the neural operator field.
Evaluation Metrics. We train our model with mean squared error (MSE) metric. We further use MSE, mean absolute error (MAE), and mean absolute percentage error (MAPE) as common evaluation metrics. Additionally, for the SEVIR dataset, we add the CSI index (Ayzel, Scheffer, and Heistermann 2020) as a core metric for comparison. We place the metrics descriptions in Appendix B.
Implementation Details. We implement our model using PyTorch framework and leverage the four A100-PCIE-40GB as computing support. In our paper, we generate model configurations (ours-$Ti/S/B$) by adjusting ST block numbers equal to 6,12,24. Concretely, tiny ($Ti$), small ($S$) and Base ($B$) models have $6$, $12$ and $24$ ST Blocks, respectively. Specifically, ST block can be divided into three sub-blocks, namely spatial encoder FoTF, temporal block TeDev and decoder block. We have placed the detailed descriptions in table 4 in Appendix C.
@@
${\cal Q}1$: Scalability analysis. In our implementation, we can quickly expand the model size by stacking blocks layers. As shown in Tab 2, we further explore the scalability of the temporal and spatial module. As meteorological exhibit highly nonlinear and chaotic characteristics, we selected the SEVIR to analyze the scalability of temporal component. We conduct experiments by selecting $2\sim 14$ TeDev blocks layers, under settings with batch size as 16, training epochs as 300, and learning rate as 0.01 (Adam optimizer). The visualization results are presented quantitatively in the Fig 5. With the increase of the number of TeDev Blocks, we can observe a gradual decrease in the MSE index, and we can easily find that loacl details are becoming clearer as time module gradually increases. This phenomenon once again demonstrates the size scalability of our model. Notably, our model incorporates a ViT-like architecture, offering intriguing possibilities for future expansion in a specific context.
@@
${\cal Q}2$: Efficiency analysis. Due to the inherent complexity of solving PDEs, we select the 2DSWE dataset with PDE property as the benchmark for validating the efficiency of our model. As shown in Fig 5, we can list observation: Obs 1. EarthFarseer presents a lower training error during the whole training process. Obs 2. Our model can achieve better convergence in faster training time. Specifically, we can save nearly 3/4 of the training time compared to VideoODE, which further verify the efficiency of EarthFarseer.
${\cal Q}3$: Predicting future frames with flexible lengths. EarthFarseer can address the issues of accumulated error and delay effects in RNN-based models for predicting frames of arbitrary length. With a two-stage design, our approach restores the feature map to the input dimension, effectively preserving spatial feature information. Additionally, it utilizes a linear projection layer to expand the time channel, allowing for convenient adjustment of the output frame length. Evaluation on TaxiBJ+, PD and 2DSWE datasets reveals that EarthFarseer exhibits remarkable quantitative performance in experiments involving 10 $\to$ 30, 10 $\to$ 60 and 50 $\to$ 50 frames (Fig 10, 6 and 14). We also showcase the 20 $\to$ 80 frames results on 2DSWE with different backbones, EarthFarseer outperforms baseline models the large margins (Tab 5), highlighting EarthFarseer’s exceptional flexibility and prediction accuracy. These findings position EarthFarseer as a promising method in ST domains.
Figure 6: Visualization of inflow and outflow prediction results on the TaxiBJ+ dataset. Aigned with Fig 9 in Appendix D.
${\cal Q}4$: Local fidelity analysis.
@@
We proceed to consider another issue, i.e., local fidelity problem. We choose TaxiBJ+, SEVIR and MovingMNIST dataset as validation datasets. As shown in Fig 7, our findings indicate that EarthFarseer effectively maintains local details while preserving the overall global context, particularly in the case of local outliers. This ability allows for the achievement of high fidelity in preserving local information. These results substantiate the local awareness exhibited by our model. For ease of understanding, we present the more complete visualizations in Appendix D (Fig 11 $\sim$ 12).
@@
In this part, we further explore the effectiveness of each individual component. In our settings, A denotes remove local convolutional component and D represents replace decoder with linear convolutional decoder. As shown in Tab 3, our ablation experiments demonstrate that removing any module from our model leads to varying degrees of performance degradation. Both the local constraint (LC) and global constraint (FoTF) contribute to the model’s performance on the spatial modules. For example, on the TaxiBJ+ dataset, removing the LC module leads to a decrease of 0.7 in the MAE metric, and removing the FoTF module leads to a decrease of 0.3. Our TeDev module outperforms the models using ViT and SwinT as a replacement for TeDev on the MovingMNIST and PD datasets. This suggests that our TeDev module is more suitable for tasks involving temporal information than using Transformer models. In summary, our model exhibits stronger expressive power in all scenarios, demonstrating the effectiveness of our proposal.
@@ Conclusion
This paper addresses the limitations of existing models that arise from the meticulous reconciliation of various advantages. We conducted a systematic study on the shortcomings faced by such models, including low scalability, inefficiency, poor long output predictions, and lack of local fidelity. We propose a scalable framework that combines spatial local-global information extraction module and temporal dynamic evolution module. EarthFarseer demonstrates strong adaptability across various tasks and datasets, exhibiting fast convergence and high local fidelity in long-distance prediction tasks. Through extensive experiments and visualizations conducted on eight physical datasets, we showcase the SOTA performance of our proposal.
Acknowledgement

### v3

@@ Introduction
Modeling spatio-temporal (ST) physical dynamics involves estimating states and physical parameters from a sequence of observations (Benacerraf 1973; Newell 1980). Generally, the understanding of a physical process is based on plenty of physical laws, such as Newton’s second law (Pierson 1993) and Conservation of energy law (Sharan et al. 1996; Egan and Mahoney 1972). As tailor-made techniques, dynamical systems, primarily rooted in diverse physical systems, have been demonstrated to conform to most fundamental principles of real-world physical phenomena (Chmiela et al. 2017; Greydanus, Dzamba, and Yosinski 2019), where such phenomena can be mostly recognized by existing mathematical frameworks. To this end, the modeling of dynamic systems has increasingly become a generic approach that yields numerous versatile techniques for various applications. It is crucial in climate science for predicting environmental impacts, in economics and finance for market analysis, in engineering for system design, in public health for disease outbreak modeling, in neuroscience for understanding brain activity, in robotics for AI development, and in logistics for optimizing supply chains. (Hale and Koçak 2012; Humar 2012). Actually, such dynamic systems can naturally model the time-varying evolution including both intricate natural system like meterology dynamics (Wiggins, Wiggins, and Golubitsky 2003; Humar 2012; Harish and Kumar 2016) and complex society system of human mobility like traffic evolutions (Ji et al. 2022; Chen et al. 2022a).
@@
Intriguingly, the fascination of machine learning with physical phenomena has significantly intensified in recent years (Shi et al. 2015; Greydanus, Dzamba, and Yosinski 2019). Many research efforts have turned to video understanding (Chang et al. 2021; Gao et al. 2022b) and physical-guided (Jia et al. 2021; Lu et al. 2021) deep learning to capture ST characteristics in a data-driven manner. These approaches usually design various spatial or temporal components tailored for effectively characterizing a specific scenario or a dedicated task. Although many ST frameworks deliver higher accuracy than the simple ones, they inevitably suffer from partial drawbacks ($\cal D$) outlined below:
${\cal D}1$: Lack of local fidelity. Given the inconsistency and distinctive local-global dependencies in dynamic systems, existing modules usually focus on global regularity but fail to preserve the local fidelity – local dynamics may differ from global dynamics.
${\cal D}2$: Poor predictions over long time-steps. Complex and continuous dynamic systems often exhibit intricate temporal correlations, leading to poor performance in long-term predictions (Isomura and Toyoizumi 2021).
${\cal D}3$: Low scalability. The intricate and convoluted component designs confine the model only capable to resolve specific tasks (e.g., super-resolution (Liang et al. 2019), flow prediction (Pan et al. 2019b)), leading to limited scalability.
${\cal D}4$: Inefficiency. The well-designed but cumbersome ST blocks not only contribute to the inefficient training (Wang et al. 2023a, 2022a) process but also pose challenges to the model deployment.
@@
For temporal correlations, we design a temporal dynamic evolution module, TeDev, which effectively captures the continuous dynamic evolution within low-dimension space. In comparison to traditional modeling of discrete static frames (Walker, Razavi, and Oord 2021; Wang et al. 2022c), our model undergoes a transformation from the continuous time domain to the frequency domain through Fourier transformation, better preserving the long-term dependence of spatio-temporal data. Further, different from the prevailing neural ordinary differential equation (ODE) algorithms (Park et al. 2021), our model showcases the great prominence in efficiency and long-term prediction tasks when captures intricated dynamics of systems without differential equation based nonlinear features. Through the exploitation of a low-parameter linear convolution projection, TeDev can effciently and accurately predict arbitrary future frames (right side of Fig 2).
@@
Spatio-temporal prediction methods
can be roughly divided into CNN-based (Oh et al. 2015; Mathieu, Couprie, and LeCun 2015; Tulyakov et al. 2018), RNN-based (Wang et al. 2023b, 2022c), and other models including the combinations (Weissenborn, Täckström, and Uszkoreit 2019; Kumar et al. 2019) and transformer based models (Dosovitskiy et al. 2020; Bai et al. 2022). While there are several existing models based on graph neural networks (GNNs), their primary focus is on handling graph data (Wang, Cao, and Philip 2020; Wang et al. 2022b), which go out of the scope of our work.
Video Prediction
has become a crucial research topic in the multimedia community, resulting in the proposal of numerous methods to tackle this challenge. Early studies primarily focused on analyzing spatio-temporal signals extracted from RGB frames (Shi et al. 2015). Recently, there has been a growing interest in integrating video prediction with external information such as optical flow, semantic maps, and human posture data (Liu et al. 2017; Pan et al. 2019a). However, in real-world applications, accessing such external information may not always be feasible (Wu et al. 2023; Anonymous 2023). Moreover, the current solutions still exhibit suboptimal efficiency and effectiveness when dealing with high-resolution videos. In this study, we concentrate on modeling continuous physical observations, which can sometimes be interpreted as a video prediction task.
@@
Dynamical systems describe how a system evolves from its current state to future states. Specifically, with an environment space $X$ and a state space $S\subset X$, they’re modeled by a function $F$.
@@
Fig 3(b) illustrates the stream morph operator converting the discrete ST sequence into a continuous stream, processed through the MSFC module and Fourier-based unit for hidden feature extraction. This operator merges channel and time dimensions to transform discrete temporal dynamics into a continuous, irregular shape. The MSFC module leverages a multi-scale fully convolutional architecture with different kernel sizes for broad and detailed feature extraction—larger kernels capture global patterns, while smaller ones focus on local details. TeDev also includes an FFT/IFFT module (see Eq $4\sim 6$) for transforming signals between time and frequency domains, enhancing signal analysis. TeDev effectively captures information across multiple time scales, ensuring temporal detail preservation in spatiotemporal predictions through its integrated approach.
@@
In this section, we empirically demonstrate the superiority of our framework on seven datasets, including two human social dynamics system (TaxiBJ+, KTH (Schuldt, Laptev, and Caputo 2004)), five natural Scene dynamical systems (SEVIR (Veillette, Samsi, and Mattioli 2020), RainNet (Chen et al. 2022b), Pollutant-Diffusion (PD), Reaction-Diffusion) and 2D shallow water Equations (2DSWE) (Takamoto et al. 2022), and a synthetic systems (MovingMNIST (Srivastava, Mansimov, and Salakhudinov 2015)). In the subsequent section, we will provide a detailed introduction to the dataset and baseline, along with the corresponding experimental settings and results.
@@
${{\cal B}_{1}}$. Video Prediction Models: We select ConvLSTM (Shi et al. 2015), PredRNN-v2 (Wang et al. 2022c), E3D-LSTM (Wang et al. 2019) and SimVP-v2 (Tan et al. 2022) as some of the most representative and advanced RNN architecture models in recent years.
${{\cal B}_{2}}$. Spatio-temporal Series Modeling: We conduct experiments on advanced Transformer architecture models, including Vision Transformer (Dosovitskiy et al. 2020), Swin Transformer (Liu et al. 2021), Rainformer (Bai et al. 2022) and Earthformer (Gao et al. 2022a).
${{\cal B}_{3}}$.Physics-guided Neural Networks: We use modeling methods that incorporate PDE or ODE in the model as baseline models for comparison, including Ad-fusion model, PhyDnet (Guen and Thome 2020), Vid-ODE (Park et al. 2021), PDE-STD (Donà et al. 2020), and FourcastNet (Pathak et al. 2022), and FourcastNet represents a meteorological modeling model in the neural operator field.
Evaluation Metrics
We train our model with mean squared error (MSE) metric. We further use MSE, mean absolute error (MAE), and mean absolute percentage error (MAPE) as common evaluation metrics. Additionally, for the SEVIR dataset, we add the CSI index (Ayzel, Scheffer, and Heistermann 2020) as a core metric for comparison. We place the metrics descriptions in Appendix B.
Implementation Details
We implement our model using PyTorch framework and leverage the four A100-PCIE-40GB as computing support. In our paper, we generate model configurations (ours-$Ti/S/B$) by adjusting ST block numbers equal to 6,12,24. Concretely, tiny ($Ti$), small ($S$) and Base ($B$) models have $6$, $12$ and $24$ ST Blocks, respectively. Specifically, ST block can be divided into three sub-blocks, namely spatial encoder FoTF, temporal block TeDev and decoder block.
@@
${\cal Q}1$: Scalability analysis
In our implementation, we can quickly expand the model size by stacking blocks layers. As shown in Tab 2, we further explore the scalability of the temporal and spatial module. As meteorological exhibit highly nonlinear and chaotic characteristics, we selected the SEVIR to analyze the scalability of temporal component. We conduct experiments by selecting $2\sim 14$ TeDev blocks layers, under settings with batch size as 16, training epochs as 300, and learning rate as 0.01 (Adam optimizer). The visualization results are presented quantitatively in the Fig 5. With the increase of the number of TeDev Blocks, we can observe a gradual decrease in the MSE index, and we can easily find that loacl details are becoming clearer as time module gradually increases. This phenomenon once again demonstrates the size scalability of our model.
@@
${\cal Q}2$: Efficiency analysis
Due to the inherent complexity of solving PDEs, we select the 2DSWE dataset with PDE property as the benchmark for validating the efficiency of our model. As shown in Fig 5, we can list observation: Obs 1. EarthFarseer presents a lower training error during the whole training process. Obs 2. Our model can achieve better convergence in faster training time. Specifically, we can save nearly 3/4 of the training time compared to VideoODE, which further verify the efficiency of EarthFarseer.
${\cal Q}3$: Predicting future frames with flexible lengths
EarthFarseer can address the issues of accumulated error and delay effects in RNN-based models for predicting frames of arbitrary length. With a two-stage design, our approach restores the feature map to the input dimension, effectively preserving spatial feature information. Additionally, it utilizes a linear projection layer to expand the time channel, allowing for convenient adjustment of the output frame length. Evaluation on TaxiBJ+, PD and 2DSWE datasets reveals that EarthFarseer exhibits remarkable quantitative performance in experiments involving 10 $\to$ 30, 10 $\to$ 60 and 50 $\to$ 50 frames. We also showcase the 20 $\to$ 80 frames results on 2DSWE with different backbones, EarthFarseer outperforms baseline models the large margins, highlighting EarthFarseer’s exceptional flexibility and prediction accuracy. These findings position EarthFarseer as a promising method in ST domains.
Figure 6: Visualization of inflow and outflow prediction results on the TaxiBJ+ dataset.
${\cal Q}4$: Local fidelity analysis
@@
We proceed to consider another issue, i.e., local fidelity problem. We choose TaxiBJ+, SEVIR and MovingMNIST dataset as validation datasets. As shown in Fig 7, our findings indicate that EarthFarseer effectively maintains local details while preserving the overall global context, particularly in the case of local outliers. This ability allows for the achievement of high fidelity in preserving local information. These results substantiate the local awareness exhibited by our model.
@@
In this part, we further explore the effectiveness of each individual component. In our settings, A denotes remove local convolutional component and D represents replace decoder with linear convolutional decoder. As shown in Tab 3, our ablation experiments demonstrate that removing any module from our model leads to varying degrees of performance degradation. Both the local constraint (LC) and global constraint (FoTF) contribute to the model’s performance on the spatial modules. For example, on the TaxiBJ+ dataset, removing the LC module leads to a decrease of 0.7 in the MAE metric, and removing the FoTF module leads to a decrease of 0.3. Our TeDev module outperforms the models using ViT and SwinT as a replacement for TeDev on the MovingMNIST and PD datasets. This suggests that our TeDev module is more suitable for tasks involving temporal information than using Transformer models.
@@ Conclusion
This paper addresses the limitations of existing models that arise from the meticulous reconciliation of various advantages. We conducted a systematic study on the shortcomings faced by such models, including low scalability, inefficiency, poor long output predictions, and lack of local fidelity. We propose a scalable framework that combines spatial local-global information extraction module and temporal dynamic evolution module. EarthFarseer demonstrates strong adaptability across various tasks and datasets, exhibiting fast convergence and high local fidelity in long-distance prediction tasks. Through extensive experiments and visualizations conducted on eight physical datasets, we showcase the SOTA performance of our proposal. All in all, our Earthfarseer achieves excellent performance.
Acknowledgements